OpenStudy (anonymous):

How can you prove that n! majorizes a^n (a is some real constant) by taking the limit of (a^n)/n! as n tends to infinity *without* introducing the gamma function?

OpenStudy (anonymous):

ok i give up what does "majorize" mean? the limits as \(n\to \infty\) of \(\frac{a^n}{n!}=0\) ?

OpenStudy (anonymous):

Essentially show a^n << n! by taking that limit

OpenStudy (anonymous):

this is the series for \(e^a\) which converges for all \(a\) since it is convergent, the terms must go to zero not much a proof i guess, but certainly true

OpenStudy (anonymous):

actually i like that proof, unless you do not get to use that fact you could say that after \(n>a\) each fraction is less than one except for the finite ones at the beginnig

OpenStudy (anonymous):

I'm not entirely sure I would get away with that. Thanks for your help anyway though, I think I may have it

OpenStudy (anonymous):

the second part should work though right? after \(n>a\) you have \[\frac{a}{1}\times \frac{a}{2}...\times \frac{a}{n}\] and \(\frac{a}{n}<1\) so after that, each fraction is less than one, and there are an infinite number of them only the first ones with \(a\geq n\) are larger than one

OpenStudy (anonymous):

Yes, so I've named the product of the first a terms M. So 0 < a^n/n! < M(a/n) and M(a/n) -> 0 as n → ∞ therefore a^n/n! → 0 by the squeeze (sandwich) theorem